QUESTION 2 2.1. Find the change of basis matrix P∁←ℬ for the bases

ℬ = {(9, 2), (4, −3)} and ∁= {(2, 1), (−3, 1)} of ℝ2 .


2.2.Verify [v]∁ = P∁←ℬ[v]ℬ for v = (−5, 3). 


1
Expert's answer
2022-04-17T09:38:34-0400

2.1. Denote "a_1=(9,2),a_2=(4,-3)" and "b_1=(2,1),b_2=(-3,1)". Any vector "v" can be presented as: "v=\alpha_1a_1+\alpha_2a_2" or "v=\beta_1b_1+\beta_2b_2" with real "\alpha_i,\beta_j" . The aim is to find the matrix "T" such that: "\left(\begin{array}{ll}\beta_1\\\beta_2\end{array}\right)=T\left(\begin{array}{ll}\alpha_1\\\alpha_2\end{array}\right)". We may consider two vectors: "v_1=\alpha^1_1a_1+\alpha^1_2a_2,""v_1=\beta^1_1b_1+\beta^1_2b_2" and "v_2=\alpha^2_1a_1+\alpha^2_2a_2,v_2=\beta^2_1b_1+\beta^2_2b_2". Then consider equation: "\left(\begin{array}{ll}\beta^1_1&\beta^2_1\\\beta^1_2&\beta^2_2\end{array}\right)=T\left(\begin{array}{ll}\alpha^1_1&\alpha^2_1\\\alpha^1_2&\alpha^2_2\end{array}\right)" and receive: "\left(\begin{array}{ll}\beta^1_1&\beta^2_1\\\beta^1_2&\beta^2_2\end{array}\right)\left(\begin{array}{ll}\alpha^1_1&\alpha^2_1\\\alpha^1_2&\alpha^2_2\end{array}\right)^{-1}=T" . As an example we take: "v_1=a_1" and "v_2=a_2". Then, "\left(\begin{array}{ll}\alpha^1_1&\alpha^2_1\\\alpha^1_2&\alpha^2_2\end{array}\right)=\left(\begin{array}{ll}1&0\\0&1\end{array}\right)". "a_1=\beta^1_1b_1+\beta^1_2b_2, a_2=\beta^2_1b_1+\beta^2_2b_2" provide two systems of equations: "9=2\beta_1^1-3\beta_2^1", "2=\beta_1^1+\beta_2^1". We receive: "\beta_1^1=3,\beta_2^1=-1". Another system of equations is: "4=2\beta_1^2-3\beta_2^2", "-3=\beta_1^2+\beta_2^2" . We get: "\beta_1^2=-1,\beta_2^2=-2." Thus, matrix "T" has the form: "T=\left(\begin{array}{ll}3&-1\\-1&-2\end{array}\right)".

2.2. The aim is to find representations of vector "v" in basis "B" and basis "C". We have: "v=\beta_1b_1+\beta_2b_2". We receive equations: "2\beta_1-3\beta_2=-5,\beta_1+\beta_2=3." The solution is: "\beta_1=\frac45,\beta_2=\frac{11}{5}". From representation: "v=\alpha_1a_1+\alpha_2a_2" we get: "9\alpha_1+4\alpha_2=-5,2\alpha_1-3\alpha_2=3." We get: "\alpha_1=-\frac{3}{35},\alpha_2=-\frac{37}{35}". In order to check the transformation, multiply: "T\left(\begin{array}{ll}\alpha_1\\\alpha_2\end{array}\right)=\left(\begin{array}{ll}3&-1\\-1&-2\end{array}\right)\left(\begin{array}{ll}-\frac{3}{35}\\-\frac{37}{35}\end{array}\right)=\left(\begin{array}{ll}\frac{4}{5}\\\frac{11}{5}\end{array}\right)". Thus, the transformation is correct.